# Portable, Massively Parallel Implementation of a Material Point Method for Compressible Flows

### v2

@@ 3 Portable MPM Implementation
$\qquad m\mathbf{v}=\mathrm{P2G}(M\mathbf{V})$
9:     Boundary condition enforcement on $m\mathbf{v}$
@@
This algorithm is of special interest for acceleration on GPUs since it can be written in terms of the four kernels (15)-(18), which present interesting parallelisation potentialities. Grid-to-Particles operations are intrinsically Single Instruction Multiple Data (SIMD) parallelisable, since data is projected from grid nodes to particles, and thus particles are independent from one another during writing stage. Particles-to-Grid kernels present the same potentialities, however, if they are implemented strictly following their analytical definition, they present a data race: different particles need access in both read and write mode to the same grid nodes.
@@
In this work, we chose to patch the Particles-to-Grid and Grid-to-Particles kernels written in the external quadgrid library to adapt them to GPU execution via Thrust rather than writing ad hoc versions in the code, at the expense of some further optimisation that could have been added by exploiting the peculiarities of our test cases, to arrive at more general and portable conclusions. In this way, there is only one computational kernel per operation (15)-(18). In any case, preliminary tests have shown that applying a kernel-fusion approach [44] where possible wouldn’t have affected the execution time order of magnitude.
@@
To better illustrate the concepts introduced in the last paragraph, we sketch and discuss below the implementation and use of the two main kernels in the algorithm, namely the Grid-to-Particle (G2P) operation (13) and the Particle-to-Grid operation (12). A simplified sketch of the implementation of the former is given in Fig. 1, while the latter is described in Fig. 4; the reported code corresponds roughly to the actual implementation available in the quadgrid library except for a few details that are left out only to avoid distraction as they are not interesting for the current discussion. As already noted, the G2P operation is easily implemented via a loop over all particles and lends itself quite naturally to SIMD parallelisation as all write operations are independent and may be carried out concurrently. The actual looping is performed via the for_each method, which can be used to perform any operation on any number of variables associated to a particle by using a counting iterator and a suitably constructed callable object as shown in Fig. 3 and Fig. 2. The P2G operation is implemented along lines very similar to those followed for G2P. Specifically, as shown in Fig. 4 the looping is done over the list of particles just as for G2P, which is made possible by the index array ptcl_to_grid that lists, for each particle, the grid cell within which it is located. Efficiently building the latter data structure heavily depends on the use of structured grid. The main difference between P2G and G2P consists in the need, in the former, to work around a data race that is created when threads working on different particles try to write data related to the same neighbouring cell vertex. The simplest approach, followed in this work, is to make the write operations atomic by using atomicAdd (as shown for sake of example in Fig. 4), which is readily available in CUDA but can also be re-implemented in standard C++ based on the std::atomic_ref class template introduced in the C++20 standard. Although our tests show that the use of atomics does not seem to lead to much of a performance degradation on GPUs, we discuss in a later section a possible alternative based on data reordering and mesh colouring that can be used to avoid atomics.
@@ 4 Benchmarks
In subsection 4.3 we then test the efficiency and performance of the implementation in providing parallel scalability on NVIDIA A100 GPUs. Lastly, in subsection 4.4, we assess the portability of the developed code to different CPUs and GPUs architecture, which is our main claim.
The experiments were carried out on NVIDIA A100 SXM4 40 GB GPUs, Intel Xeon Platinum 8260 and AMD EPYC Rome 7402 CPUs at Leonardo S.p.A.’s in-house HPC facility davinci-1, and on AMD EPYC 7313, AMD EPYC 7413 CPUs and AMD Radeon Instinct MI210 GPUs at MOX Laboratory for Modeling and Scientific Computing at Politecnico di Milano’s Mathematics Department.
All tests are performed in double precision, i.e. using FP64. Depending on the architecture, better performance could be achieved with lower precision floating points; for example, the NVIDIA GPUs here employed, which date back to 2020, should gain a 2x factor in computational time when working with FP32. However, despite the code supports different choices, we have focused only on doubles, since floats could lead to numerical difficulties in some test cases, and already the slightly more recent AMD MI 210 GPUs (release date December 2021) are able to handle double and single precision floats without any performance degradation.
@@ 4.1 Test cases
In all test cases except in the Sod Shock Tube, initial particles position is sampled by a uniform random distribution in the fluid domain. Indeed, a sufficient randomness of particles position, especially in particles recycling from outlet to inlet, is fundamental to obtain stable and physically meaningful simulations, in agreement with results in the literature on Particle-In-Cell methods in plasma physics. On longer runs (i.e. above tens of physical seconds) a symplectic time integration scheme might be useful [34].
@@
In order to reproduce the 1D problem, we introduce a mesh of $1\times 100$ cells; moreover, to be able to compare the MPM solution with the reference one, we also set initial particle position as in [65, 64].
@@
Supersonic flow past solid obstacles
We then consider two cases of supersonic, Mach 3 flow:
  • (a)
    past a cylinder of radius $0.25$ and centre $(0.6,1)$ in a $[0,4]\times[0,2]$ rectangular domain; and
  • (b)
    past a rectangular-shaped step $[0.6,3]\times[0,0.2]$ in a $[0,3]\times[0,1]$ rectangular domain.
In both cases, we enforce wind tunnel–like boundary conditions, with inflow conditions at the left vertical boundary, outflow conditions at the right horizontal boundary, and slip conditions at the top and bottom boundaries.
@@
$\begin{cases}\mathbf{v}_{inflow}=(3,0)\\ \mathbf{v}_{top}\cdot\mathbf{n}=\mathbf{v}_{bottom}\cdot\mathbf{n}=0\end{cases}$
where $\mathbf{n}$ denotes the unit normal vector. A particles-recycling approach is adopted, whereby particles leaving the domain at the outflow boundary are regenerated in a random cell at the inflow boundary with inflow conditions, keeping the total number of particles constant throughout the simulation. In test case (a), as gas particles travel past the cylinder, a shock wave develops at the front and two oblique shocks in the rear. The first shock is reflected by walls and interacts with the back shocks; see Figure 6 and its caption for numerical details.
@@
Figure 7: Mach 3 flow past a step: pressure at $t=12s$, $N_{C}=326.7\cdot 10^{3}$, $ppc=16$, $N_{P}=4.4\cdot 10^{6}$
In the step test (b), a frontal shock develops that is later reflected from the top wall and successively from the step see Figure 7 and its caption for numerical details.
The MPM implementation is able to correctly simulate the main flow features, in case of both round shapes and sharp edges and corners. Computed shock-waves angle and flow deflection angle (Figure 6) are found to be in agreement with experimental references [20] as well as with results with other codes, e.g., deal.II, when the same polynomial order is used for basis functions. In particular, the cylinder test case is simulated in the deal.II documentation [39] [^7] https://www.dealii.org/current/doxygen/deal.II/step_69.html#Results. With our implementation, we are able to reproduce the main features of the front shock starting from a very low resolution, with details, especially in the rear, getting better as resolution increases; more details are shown in subsection 4.2. Overall, the results are in agreement with [39], while the lower level of detail in comparison with some results in the literature ([16, 38, 27]) is likely due to the first order space discretisation used here.
@@
Taylor-Green vortex
In the third test, we address the Taylor-Green Vortex test case, a well studied, fluid-only benchmark to assess both viscous and inviscid fluid dynamics codes in the literature, usually at subsonic regime ([11, 7, 54]); here we simulate the case in supersonic Mach 3 conditions.
In the inviscid case, the dynamics is strongly affected by problem symmetries, leading to “impermeable box boundary conditions” [7]. These symmetries are:
  • within the $[0,\pi]^{2}$ square-shaped computational domain, the fluid never crosses the lines $x=n\pi$ and $y=n\pi$; and
  • close to the boundaries, both the normal velocity $\mathbf{v_{\perp}}$ and the normal derivative ${\tfrac{\partial\mathbf{v_{//}}}{\partial\mathbf{n}}}$ of the tangential velocity $\mathbf{v_{//}}$ vanish.
We then set the initial conditions:
$\begin{cases}v_{x}(x,y)=v_{\infty}\sin\left(\tfrac{x}{\pi}\right)\cos\left(\tfrac{y}{\pi}\right)\\ v_{y}(x,y)=-v_{\infty}\cos\left(\tfrac{x}{\pi}\right)\sin\left(\tfrac{y}{\pi}\right)\\ p(x,y)=p_{\infty}+\tfrac{\rho_{\infty}}{16}\left(\left(\cos\left(\tfrac{2x}{\pi}\right)+\cos\left(\tfrac{2y}{\pi}\right)\right)\cdot 3-2\right).\end{cases}$
The numerical method captures the flow features with good accuracy as compared to results in the literature (Figure 10, cf. Figure 1b in [7]), confirming the general applicability of the MPM algorithm in simulating compressible fluid flow.
Figure 10: Taylor-Green vortex: computed velocity at $t=2s$, $N_{C}=3.94\cdot 10^{5}$, $ppc=9$, $N_{P}=3.55\cdot 10^{6}$
@@ 4.2 Sensitivity to numerical parameters
Besides resolution, supersonic compressible fluid dynamics requires further care in setting the average number of particles per cell. Indeed, $ppc$ affects stability in a twofold fashion.
First, for low $ppc$ values (e.g.,1), widespread oscillations are observed all along the simulation, due to almost randomly distributed empty cells. In a recent study, a $ppc$ value of 4 was found to be optimal for convergence of gas dynamics tests at lower speed [50], and, indeed, a $ppc=4$ solves this issue. However, for supersonic flow past a static obstacle, we observed that using $ppc=4$ might cause the formation of domain regions of empty or quasi-empty cells, which gets evident as grid resolution increases, during the initial transient phase from uniform, undisturbed conditions, as can be seen from Figure 13. Such emptying causes later formation of nonphysical cavitation-like effects during the transient that precedes equilibrium, which alter the dynamics, and, if particularly strong, can cause also spurious shock waves, that interact with the physical ones. These events can be prevented by increasing the average particles per cell number: at Mach 3, spurious shock waves become less relevant with $ppc=9$, and undetectable with $ppc=16$.
@@ 4.3 Computational performance
We then focus on the computational performance and scalability of the method, analysing the computational time required by simulations when varying the grid resolution and $ppc$ values. It is worth noting that the implemented parallel kernels can be particle- or grid-based, and, respectively, assign one particle or one grid node to one GPU thread. Moreover, we implemented both mapping kernels (P2G and G2P) as particle-based kernels, since this proved to be the most efficient option.
@@
The first finding is that both kernels scale linearly with the total number of particles, in agreement with their implementation (Figures 15 and 16), while pure grid kernels, such as boundary conditions enforcement, scale with the number of nodes. For this reason, we applied a global particles reordering step in the algorithm and studied its effects on performance. Indeed, the ordering is done so that particles which are in the first grid cell after moving are assigned the lowest global indices, followed by those in the second grid cell, and so on moving along the grid. Thus, data which will be accessed at the next time iteration will be ordered, with contiguous particles having contiguous indexes, keeping data locality across the simulation, which is of utmost importance on our target architecture. The reordering turns out to have a significant impact on the G2P kernels, resulting in a further 4x speed-up factor, while preserving the linear scaling (Figure 15).
@@
By contrast, the reordering appears not to have an appreciable effect on the P2G kernels, with slowdowns and accelerations evident at different particle values (Figure 16). The reason can be found in the main implementation difference among the two kernels. P2G introduces a data race, that is currently dealt with using CUDA’s atomicAdd, and its cost becomes more apparent with re-ordering. Indeed, even though, to the best of the authors’ knowledge, NVIDIA’s A100 scheduler policy is not accessible, numerical examples appear to suggest that the scheduler behaves more in a round-robin fashion rather than in a block partitioning way. So, a particle is assigned to a GPU thread, until there are some free, then a second particle is assigned, and so on. In this framework, randomisation of the global order of particles is somewhat beneficial, since it lowers the probability that particles that are in the same cells are processed simultaneously. After reordering, particles close in the domain are also close in memory, enhancing probability of a memory collision, i.e. of simultaneous access to the same data. Since in our case such access would be in read-and-write mode, it activates the atomic operation and degrades performance.
@@
This also explains why reordering appears to not being beneficial for P2G kernels. Since a higher number of ordered particles per cell increases the chance of data races, the average number of particles per cell becomes an important parameter. Indeed, grouping simulations results per $ppc$ value, we recover linearity, and can also confirm the trend of worsening performance at larger numbers of ordered particles in a cell. Indeed, comparing the linear fits in Figure 17, we can see that in the $ppc=4$ and $ppc=9$ cases execution is faster than in the global fit, and in the $ppc=16$ case it is slower.
@@
noindent As for overall computational time, we note that its value largely depends on the requested resolution. On a NVIDIA A100 GPU, to simulate one second of gas flow around the cylinder with $4\cdot 10^{4}$ particles requires about one second of computational time on the device, with $16\cdot 10^{6}$ particles it requires about $3$ minutes. This is reasonably fast, especially considering that the latter resolution is rather excessive for the 2D cylinder test case. For comparison, 3D snow avalanches were recently simulated via MPM with $10$ to $30$ million particles, see [25].
@@
In the next paragraph, we exploit the flexibility granted from this approach to test the code on different hardware architectures, both on CPUs and GPUs.
@@ 4.4 Portability
For the first test, we vary the number of cells and of particles per cell and study how this affects the main computational kernels on an AMD CPU, employing $242x$ 24 cores in parallel (Figure 18).
While portability is automatically granted, CPU performance turns out to be worse than single-GPU performance, even though $242x$ 24 CPU cores are employed in parallel. This is consistent with expectations that MPM is better suited to GPU architectures.
@@
For the second test, we focus on the heavier test case, the cylinder with $16$ million particles, and simulate one physical second on different architectures, comparing the total wall clock time, referred to the time required by the device, being either a CPU or a GPU; see Figure 19. It is interesting to note that the chosen implementation approach allows the porting of the code to AMD GPUs without any modification.
@@
Lastly, we focus on the comparison between the two GPU architectures.
@@
However, at heavier loads, despite the comparable specs, the AMD MI210 starts to display longer wallclock times than the NVIDIA A100, see Figure 21 for a comparison of the overall computational time on the device. This difference might be due to different aspects, such as having developed the code, and taken performance based design choices, only on NVIDIA A100 GPUs, and having ported it to different hardware afterwards. However, GPU performance is overall competitive with respect to CPU. For example, the heavier test case is 7 to 8 times slower on MI210 than on A100, but anyway 3 to 4 times faster than parallel execution via Intel OneAPI-TBB on Intel Xeon with 2x24 cores.
@@
Figure 20: Comparison of performance on different GPUs: analysis of the average computational time for a low resolution test case $N_{C}=10^{4}$, $ppc=4$, corresponding to about half computational second to simulate 1 physical second.

### v4

@@ 3 Portable MPM Implementation
$\qquad\mathbf{mv}=\mathrm{P2G}(\mathbf{MV})$
9:     Boundary condition enforcement on $\mathbf{mv}$
@@
This algorithm is of special interest for acceleration on GPUs since it can be written in terms of the four kernels (16)-(19), which present interesting parallelisation potentialities. Grid-to-Particles operations are intrinsically Single Instruction Multiple Data (SIMD) parallelisable, since data is mapped from grid nodes to particles, and thus particles are independent from one another during writing stage. Particles-to-Grid kernels present the same potentialities, however, if they are implemented strictly following their analytical definition, they present a data race: different particles need access in both read and write mode to the same grid nodes.
@@
In this work, we chose to patch the Particles-to-Grid and Grid-to-Particles kernels written in the external quadgrid library to adapt them to GPU execution via Thrust rather than writing ad hoc versions in the code, at the expense of some further optimisation, such as specific kernel-fusions, that could have been added by exploiting the peculiarities of our model (5), to arrive at more general and portable conclusions. This way, there is only one computational kernel per operation (16)-(19). In any case, preliminary tests have shown that applying a kernel-fusion approach [47] where possible wouldn’t have affected the execution time order of magnitude.
@@
To better illustrate the concepts introduced in the last paragraph, we sketch and discuss below the implementation and use of the two main kernels in the algorithm, namely the Grid-to-Particle (G2P) operation (17) and the Particle-to-Grid operation (16). A simplified sketch of the implementation of the former is given in Fig. 1, while the latter is described in Fig. 4; the reported code corresponds roughly to the actual implementation available in the quadgrid library except for a few details that are left out only to avoid distraction, as they are not interesting for the current discussion. As already noted, the G2P operation is easily implemented via a loop over all particles and lends itself quite naturally to SIMD parallelisation as all write operations are independent and may be carried out concurrently. The actual looping is performed via the for_each method, which can be used to perform any operation on any number of variables associated to a particle by using a counting iterator and a suitably constructed callable object as shown in Fig. 3 and Fig. 2. The P2G operation is implemented along lines very similar to those followed for G2P. Specifically, as shown in Fig. 4 the looping is done over the list of particles just as for G2P, which is made possible by the index array ptcl_to_grid that lists, for each particle, the grid cell within which it is located. Efficiently building the latter data structure heavily depends on the use of structured grid. The main difference between P2G and G2P consists in the need, in the former, to work around a data race that is created when threads working on different particles try to write data related to the same neighbouring cell vertex. The simplest approach, followed in this work, is to make the write operations atomic by using atomicAdd (as shown for sake of example in Fig. 4), which is readily available in CUDA and HIP but can also be re-implemented in standard C++ based on the std::atomic_ref class template introduced in the C++20 standard. Although our tests show that the use of atomics does not seem to lead to much of a performance degradation on GPUs, we discuss in a later section a possible alternative based on data reordering and mesh colouring that can be used to avoid atomics.
@@ 4 Benchmarks
In subsection 4.3 we then test the efficiency and performance of the implementation in providing parallel scalability on NVIDIA A100 GPUs. Lastly, in subsection 4.4, we assess the performance portability of the developed code to different CPUs and GPUs architecture, which is our main claim.
The HPC experiments were carried out on NVIDIA A100 SXM4 40 GB GPUs, Intel Xeon Platinum 8260 and AMD EPYC Rome 7402 CPUs at Leonardo S.p.A.’s in-house HPC facility davinci-1, and on AMD EPYC 7313, AMD EPYC 7413 CPUs and AMD Radeon Instinct MI210 GPUs at MOX Laboratory for Modeling and Scientific Computing at Politecnico di Milano’s Mathematics Department.
All tests are performed in double precision, i.e. using FP64. Depending on the architecture, better performance could be achieved with lower precision floating points; for example, the NVIDIA A100 GPUs employed here should gain a 2×2$\times$2 × factor in computational time when working with FP32. However, although the code supports different choices, we have focused on doubles only, since floats could lead to numerical difficulties in some test cases, and since other GPUs, such as the AMD MI 210, are able to handle double and single precision floats without any theoretical performance degradation.
@@ 4.1 Test cases
In all test cases except in the Sod Shock Tube, initial particles position is sampled by a uniform random distribution in the fluid domain. Indeed, a sufficient randomness of particles position, especially in particles recycling from outlet to inlet, is fundamental to obtain stable and physically meaningful simulations, in agreement with results in the literature on Particle-In-Cell methods in plasma physics [36]. On longer runs (i.e. above tens of physical seconds) a symplectic time integration scheme might be useful [36, 5].
@@
At time $t=0.143$ we stop the simulation, thus particles do not leave the domain. In order to reproduce the 1D problem, we introduce a mesh of $1\times 100$ cells; moreover, to be able to compare the MPM solution with the reference one, we also set initial particle position as in [70, 69].
@@
Supersonic flow past a cylinder
We then consider the case of supersonic, Mach 3 flow past a cylinder of radius $0.25$ and centre $(0.6,1)$ in a $[0,4]\times[0,2]$ rectangular domain. We enforce wind tunnel–like boundary conditions, with inflow conditions at the left vertical boundary, outflow conditions at the right vertical boundary, and slip conditions at the top and bottom boundaries, as in [27].
@@
$\begin{cases}\mathbf{v}_{inflow}=(3,0)\\ \mathbf{v}_{top}\cdot\mathbf{n}=\mathbf{v}_{bottom}\cdot\mathbf{n}=0\end{cases}$ (30)
where $\mathbf{n}$ denotes the unit normal vector. A particles-recycling approach is adopted, whereby particles leaving the domain at the outflow boundary are regenerated in a random cell at the inflow boundary with inflow conditions, keeping the total number of particles constant throughout the simulation. In this test case, as gas particles travel past the cylinder, a shock wave develops at the front and two oblique shocks in the rear. The first shock is reflected by walls and interacts with the back shocks; see Figure 6 and its caption for numerical details.
@@
The MPM implementation is able to correctly simulate the main flow features; computed shock-waves angle and flow deflection angle (Figure 6) are found to be in agreement with experimental references [20] as well as with results which can be obtained with other codes, e.g., deal.II, when the same polynomial order is used for basis functions. In particular, the cylinder test case is simulated in the deal.II documentation [42] [^10] https://www.dealii.org/current/doxygen/deal.II/step_69.html#Results. With our implementation, we are able to reproduce the main features of the front shock starting from a very low resolution, with details, especially in the rear, getting better as resolution increases; more details are shown in subsection 4.2. Overall, the results are in agreement with [42], while the lower level of detail in comparison with some results in the literature is likely due to the first order space discretisation used here ([15, 41, 27]).
@@ 4.2 Sensitivity to numerical parameters
Besides grid resolution, supersonic compressible fluid dynamics requires further care in setting the average number of particles per cell. Indeed, $ppc$ affects stability in a twofold fashion.
First, for low $ppc$ values (e.g.,1), widespread oscillations are observed all along the simulation, due to almost randomly distributed empty cells. In a recent study, a $ppc$ value of 4 was found to be optimal for convergence of gas dynamics tests at lower speed [56], and, indeed, a $ppc=4$ solves this issue. However, for supersonic flow past a static obstacle, we observed that using $ppc=4$ might cause the formation of domain regions of empty or quasi-empty cells, which gets evident as grid resolution increases, during the initial transient phase from uniform, undisturbed conditions, as can be seen from Figure 11. Such emptying causes later formation of non-physical cavitation-like effects during the transient that precedes equilibrium, which alter the dynamics, and, if particularly strong, can cause also spurious shock waves, that interact with the physical ones. These events can be prevented by increasing the average particles per cell number: at Mach 3, spurious shock waves become less relevant with $ppc=9$, and undetectable with $ppc=16$.
@@ 4.3 Computational performance
We then focus on the computational performance and scalability of the method, analysing the computational time required by simulations when varying the grid resolution and $ppc$ values. It is worth noting that the implemented parallel kernels can be particle-based or grid-based, and, respectively, assign one particle or one grid node to one GPU thread. Moreover, we implemented both mapping kernels (P2G and G2P) as particle-based kernels, since this proved to be the most efficient option.
@@
The first finding is that both kernels scale linearly with the total number of particles, in agreement with their implementation (Figures 13 and 14), while pure grid kernels, such as boundary conditions enforcement, scale with the number of nodes. For this reason, we applied a global particles reordering step in the algorithm and studied its effects on performance. Indeed, the ordering is done so that the lowest global indices are assigned to the particles which are in the first grid cell after moving, followed by those in the second grid cell, and so on, moving along the grid. Thus, data which will be accessed at the next time iteration will be ordered, with contiguous particles having contiguous indexes, keeping data locality across the simulation, which is of utmost importance on our target architecture. The reordering turns out to have a significant impact on the G2P kernels, resulting in a further 4x speed-up factor, while preserving the linear scaling (Figure 13).
@@
By contrast, the reordering appears to have a slightly negative effect on the P2G kernels, with slowdowns which are more evident at some particle values (Figure 14). The reason can be found in the main implementation difference among the two kernels. P2G introduces a data race, that, on NVIDIA GPUs, is currently dealt with using CUDA atomicAdd, and its cost becomes more apparent with re-ordering. Indeed, even though, to the best of the authors’ knowledge, NVIDIA’s A100 scheduler policy is not accessible, numerical examples appear to suggest that the scheduler behaves more in a round-robin fashion rather than in a block partitioning way. So, a particle is assigned to a GPU thread, until there are some free, then a second particle is assigned, and so on. In this framework, randomisation of the global order of particles is somewhat beneficial, since it lowers the probability that particles that are in the same cells are processed simultaneously. After reordering, particles close in the domain are also close in memory, enhancing probability of a memory collision, i.e. of simultaneous access to the same data. Since in our case such access would be in read-and-write mode, it activates the atomic operation and degrades performance.
@@
Moreover, since a higher number of ordered particles per cell increases the chance of data races, the average number of particles per cell becomes an important parameter. As a matter of fact, grouping simulations results per $ppc$ value, we can confirm the trend of worsening performance at larger numbers of ordered particles in a cell. Indeed, comparing the linear fits in Figure 15, we can see that in the $ppc=4$ case execution is faster than in the global fit, in the $ppc=16$ case it is slower, the $ppc=9$ being intermediate between the two. Moreover, we can see that at higher $ppc$ the computational time required by the P2G kernels deviates from the linear trend. The reason can be found in the relation between the average number of particles per cell and the number of grid nodes per cell. Indeed, as the $ppc$ value grows and the number of nodes per cell stays fixed, data locality is lost sooner in the simulation: in each cell, a higher amount of particles indexes become increasingly different from the grid indexes, leading to an increase in the beneficial effect of re-ordering on parallel performance.
@@
As for overall computational time, we note that its value largely depends on the requested resolution. On a NVIDIA A100 GPU, to simulate one second of gas flow around the cylinder with $4\cdot 10^{4}$ particles requires about one second of computational time on the device, with $16\cdot 10^{6}$ particles it requires about $3$ minutes. This is reasonably fast, especially considering that the latter resolution is rather excessive for the 2D cylinder test case. For comparison, 3D snow avalanches were recently simulated via MPM with $10$ to $30$ million particles, see [25]; further considerations on extension to 3D are presented in section 5.
@@
In the next paragraph, we exploit the flexibility granted from this approach to test the code on different hardware architectures, both on CPUs and GPUs, and provide an evaluation of the achieved speed-up, with respect to serial and CPU-parallel versions.
@@ 4.4 Performance portability
For the first test, we vary the number of cells and particles per cell and study how this affects the main computational kernels on an AMD CPU, employing $242x$ 24 cores in parallel (Figure 16).
While portability is automatically granted, CPU performance turns out to be worse than single-GPU performance, even though $242x$ 24 CPU cores are employed in parallel, as can be seen comparing total wall-clock times in Figure 17; this is consistent with expectations that MPM is better suited to GPU architectures. It is anyway positive that the computational time stays linear with respect to the number of particles.
@@
Figure 17: Computational time scaling and speed-up varying number of total particles and parallel architecture.
For the second test, we focus on the heaviest test case (the cylinder with $16$ million particles) and simulate one physical second on different architectures, comparing the total wall-clock time, referred to the time required by the device, being either a CPU or a GPU; see Figure 18. It is interesting to note that the chosen implementation approach allows the porting of the code to AMD GPUs without any modification.
@@
Then, we focus on the comparison between the two GPU architectures.
@@
Instead, at heavier loads, despite the comparable specs, the AMD MI210 starts to display longer wall-clock times than the NVIDIA A100, see Figure 20 for a comparison of the overall computational time on the device. This difference might be due to different aspects, such as having developed the code, and taken performance based design choices, only on NVIDIA A100 GPUs, and having ported it to different hardware afterwards. However, in the literature, odd performance issues have been found in high-load scenarios on AMD MI100 GPUs that are currently being investigated from AMD itself [40], so we cannot exclude to be in a similar situation here.
In any case, GPU performance is overall competitive with respect to CPU. For example, the heaviest test case is 7 to 8 times slower on MI210 than on A100, but anyway 3 to 4 times faster than parallel execution via Intel OneAPI-TBB on Intel Xeon with 2x24 cores.
Lastly, it is interesting to note that, with the implemented approach, the code is portable also to non-HPC architectures, as common commercial processors, and that the wall-clock time is not excessive as long as the computational load is low enough, as can be seen in Figure 21.
@@
Figure 19: Comparison of performance on different GPUs: analysis of the total wall-clock computational time for a low resolution test case $N_{C}=10^{4}$, $ppc=4$, corresponding to less then one computational second to simulate 1 physical second.
@@
Figure 21: Comparison of performance on different commercial, non HPC architectures: analysis of the total wall-clock computational time for a low resolution test case $N_{C}=10^{4}$, $ppc=4$.